# An Agnostic Biosignature Based on Modeling Panspermia and Terraformation

### v2

@@ A.1.1 Additional Parameters
Mutation. Depending on the scenario, we include either a 0% or 10% chance of mutation for each element of the planet’s new composition, meaning that in the 10% scenario, with our length 10 composition vector, on average, 1 of the 10 elements is mutated during each terraformation event. Mutated elements are chosen from a continuous uniform random distribution of [0, 1], just like the initial compositions were chosen. Mutation is meant to represent the change in the kind of planetary observable characteristics which are compatible with life as life evolves over time.
@@ A.1.2 Flat curve
Because all simulations are initialized with planets of random compositions, the Mantel coefficient is initially 0. A flat curve shape indicates that the spread of life does not cause a significant increase to the mantel coefficient (Fig. A1). In simulations targeting the most compositionally similar planets within a positional distance cutoff, , flat curves occur when the distance cutoff is low enough that the terraformed planets run out of new target terraformation candidates (Fig. A1A bottom-most row, orange box). Flat shapes also occur in scenarios where terraformation maintains at least half of a target planet’s composition (Fig. A1B, rightmost column, orange box).
@@ A.1.3 Increasing then decreasing curve
When simulations are initialized with a single OoL, the progression of planet terraformation over time leads to an increasing then decreasing shaped Mantel coefficient curve, where the Mantel coefficient peaks after some of the planets have been terraformed, and then decreases afterwards (Fig. A1, data in green). Because the Mantel coefficient is measuring the correlation between positional and compositional distance matrices of planets, this correlation will decrease if either distance matrix becomes too homogeneous. With a single OoL, this is exactly what begins to happen with the compositional distance matrix, with the decline especially pronounced in the scenario with perfect replication (Fig. A1B, green box).

### v3

@@ A.1.1 Additional Parameters
Mutation. Depending on the scenario, we include either a 0% or 10% chance of mutation for each element of the planet’s new composition, meaning that in the 10% scenario, with our length 10 composition vector, on average, 1 of the 10 elements is mutated during each terraforming event. Mutated elements are chosen from a continuous uniform random distribution of [0, 1], just like the initial compositions were chosen. Mutation is meant to represent the change in the kind of planetary observable characteristics which are compatible with life as life evolves over time.
@@ A.1.2 Flat curve
Because all simulations are initialized with planets of random compositions, the Mantel coefficient is initially 0. A flat curve shape indicates that the spread of life does not cause a significant increase to the mantel coefficient (Fig. A1). In simulations targeting the most compositionally similar planets within a positional distance cutoff, , flat curves occur when the distance cutoff is low enough that the terraformed planets run out of new target terraforming candidates (Fig. A1A bottom-most row, orange box). Flat shapes also occur in scenarios where terraforming maintains at least half of a target planet’s composition (Fig. A1B, rightmost column, orange box).
@@ A.1.3 Increasing then decreasing curve
When simulations are initialized with a single OoL, the progression of planet terraforming over time leads to an increasing then decreasing shaped Mantel coefficient curve, where the Mantel coefficient peaks after some of the planets have been terraformed, and then decreases afterwards (Fig. A1, data in green). Because the Mantel coefficient is measuring the correlation between positional and compositional distance matrices of planets, this correlation will decrease if either distance matrix becomes too homogeneous. With a single OoL, this is exactly what begins to happen with the compositional distance matrix, with the decline especially pronounced in the scenario with perfect replication (Fig. A1B, green box).